# images/svg/

### v1

@@ V-B Recovery path refinement
The initially selected recovery path $\mathbf{P}_{\text{recovery}}$ undergoes further refinement to enhance safety and smoothness before execution. This involves three steps:
  1. 1.
    Centroid Offsetting: Points along $\mathbf{P}_{\text{recovery}}$ are pushed slightly outwards, away from the path’s geometric centroid, to increase clearance from potential obstacles near the path’s center.
  2. 2.
    Random Sampling Perturbation: Points are locally perturbed by sampling in random directions, seeking nearby collision-free states to potentially escape minor constraint violations or local minima.
  3. 3.
    Polynomial Smoothing: A third order polynomial function is fitted to segments of the path to reduce sharp turns and generate a smoother reference trajectory.
The planner then checks if $\mathbf{P}_{\text{recovery}}$ is collision-free. This refinement process is performed in a loop that continues until a collision-free path is found.
@@ VI Simulations Experiments
This section presents the simulation results for the proposed REACT method. The entire framework is implemented in C++ to ensure computational efficiency and is integrated with ROS to facilitate modularity and deployment on real-world robots. Furthermore, the open motion planning (OMPL) library [29] is utilized to compute the shortest paths using the RRT* algorithm.
@@ VI-A Simulation experimental setup
The path planner is implemented for a BlueROV2 underwater robot. An MPC approach accounts for model constraints and provides the optimal control input $\mathbf{u}^{ref}\in\mathbb{R}^{4}$, where $\mathbf{u}^{ref}=[F_{x},F_{y},F_{z},M_{z}]^{T}$ represents the forces in the $X$, $Y$, and $Z$ directions, and the moment about the $Z$-axis. The goal is to follow the desired reference trajectory $\mathbf{x}^{ref}\in\mathbb{R}^{6}$, where $\mathbf{x}^{ref}=[x_{ref},y_{ref},z_{ref},\psi_{ref}]^{T}$ contains the reference position ($x_{ref}$, $y_{ref}$, $z_{ref}$) and the reference yaw angle $\psi_{ref}$. For further details about the controller and model, the reader is referred to [30]. The entanglement-aware path planner provides $\mathbf{x}^{ref}$ in real time, ensuring the BlueROV2 avoids obstacles while following the desired trajectory.
We perform a comparative analysis of the proposed REACT method and a baseline conventional CPP (FC-Planner [8]), which does not explicitly handle entanglement. The simulation setup uses a 1/10-scale pipe structure based on the model described in [8]. The simulated onboard camera has a $70$-degree field of view. The tether constraint is implemented by specifying a maximum allowable tether length of $L_{\text{max}}=10$m. Note that the resulting planned trajectories are geometrically identical to those in the full-scale environment, simply scaled down proportionally. The primary motivation for using the scaled model is to accelerate simulation runs while preserving the validity of the results.
Coverage is computed in a manner similar to [9]. At each time step, the position and orientation of the camera are used to determine which triangles in the environment mesh are visible. A triangle is marked as visible if its centroid is within the inspection range, its surface normal faces the camera, and its projection lies within the camera’s field of view. Over time, the set of all uniquely observed triangles accumulates. The coverage at time $t$ is defined as the ratio of the number of unique visible triangles up to time $t$ to the total number of triangles in the environment.
@@ VI-B Simulations results
Table I: Coverage performance comparison between REACT and CPP baseline showing inspection time, recovery time, total mission duration, and final environmental coverage.
@@
Table II: Comparison of tether constraint compliance showing the maximum tether length reached and the total duration of tether length exceedance.
@@
Figure 4: Comparison of coverage and tether behavior over time.
@@
The performance metrics for both planners are summarized in Table I and Table II, while Figs. 4 and 5 visualize the performance of the planners in terms of environmental coverage, tether length behavior, and final trajectory configurations. The evaluation focuses on comparing mission efficiency, constraint compliance, and safety aspects between the entanglement-aware REACT method and the conventional CPP baseline approach.
@@
REACT demonstrates multiple instances of entanglement detection and resolution, as evidenced by the peaks in the tether length curve in Fig. 4(b) and the corresponding flat regions in the coverage curve in Fig. 4(a), where the ROV inspection progress stops to reroute and find paths without entanglement. This reactive replanning behavior directly results in a longer inspection time, as the system prioritizes tether safety over raw speed.
In particular, the limit on the length of the tether is rarely exceeded by REACT, as shown in Table II and Fig. 4(b). Conversely, the CPP method severely violates this constraint, reaching $31.16$m for extended periods ($327.37$s), risking unrecoverable tether entanglement. The 3D trajectory visualizations in Fig. 5 further illustrate this difference, showing the entangled tether geometry resulting from CPP versus the non-entangled configuration achieved by REACT. Therefore, REACT demonstrates superior performance for safe tethered ROV inspection.
@@ VII Real-world Experiments
This section presents real-world experimental results validating the proposed REACT method. The experiments were conducted at the underwater robotics facility of the German Research Center for Artificial Intelligence (DFKI), which features a $23$m×$\times 19$m×$\times 8$m test basin equipped with a 12-camera Qualisys motion capture system for precise state estimation.
VII-A Real-world experimental setup
The experimental platform is a BlueROV2, configured identically to the simulation setup. The test environment consists of a pipe structure with a diameter of $0.7$m and a length of $2.5$m, as shown in Fig. 6. Two planning modes, identical to those evaluated in simulation, are tested. The first is REACT, which represents the full system incorporating REACT. The second is the CPP baseline without REACT. In both modes, a helical trajectory is given as a reference around the pipe.
Figure 6: Experimental setup in the DFKI underwater basin showing the BlueROV2 and the test structure. The pipe is suspended horizontally in the water using cables.
VII-B Real-world experimental results
@@
Real-world tests illustrate the clear benefits of the REACT planner compared to the baseline CPP approach. Figure 7 shows the mission trajectories for both methods. As such, the baseline CPP planner failed to complete the inspection task. At $t=124$s, the tether entangles the pipe structure, stopping the ROV and preventing further progress. Hence, the overall coverage achieved is only $80.9\%$ (Table III).
@@
The proposed REACT planner completed the full mission, achieving $95.6\%$ coverage. The planner adapted to tether constraints throughout the mission. Unlike the simulation results, the soft tether length limit of the $8$m was exceeded in real-world tests. As such, when no entanglement-free paths to the goal are available, REACT treats the length constraints as a soft limit rather than a hard limitation. However, when entanglement is detected and there is an entanglement-free path around the pipe, REACT finds a path that detangles the ROV from the structure. This capability is demonstrated at $t=96$s when REACT detected an imminent tether entanglement. The planner computed a new path that safely untangled the ROV from the pipe structure, allowing inspection to continue. The system only exceeded the soft tether limit when necessary for mission completion while prioritizing entanglement avoidance whenever possible.
The baseline CPP planner, lacking entanglement-awareness, became permanently stuck. This demonstrates a key advantage of REACT that was not visible in the simulation study, where the physical tether forces remain unmodeled.
Although the REACT planner adds computation compared to the baseline, the maximum replanning time of 0.4822 s remained well within the onboard platform’s capabilities, allowing real-time operation. This overhead, while higher than the conventional CPP approach, was low enough to enable timely adaptation to tether entanglement events.
@@
These results demonstrate that standard planners without tether constraints fail even during relatively simple underwater 3D inspection tasks. While the baseline initially followed the planned path, it became entangled and subsequently immobilized midway through the mission. On the other hand, REACT’s ability to detect and avoid entanglement through replanning enabled a successful mission completion. These findings highlight the need for entanglement-aware planning in real underwater operations.
Table III: Coverage Performance in Real-world Trials
Planner Inspection Completion Final Coverage (%)
REACT Full 95.6
CPP Failed (Entangled) 80.9
@@
Figure 7: Real-world ROV trajectories. Without REACT, the mission fails due to tether entanglement. With REACT, the ROV completes the mission successfully. At $t=96$s, the replanner is activated, allowing REACT to disentangle the ROV and continue the inspection. In contrast, at $t=124$s, the ROV using the conventional CPP becomes stuck due to the entangled tether.
@@ VIII Conclusion
We present REACT, a real-time, entanglement-aware path planning framework for tethered underwater vehicles that supports underwater asset inspection tasks. Through comparative simulation and real-world experiments, we demonstrate that conventional planners face tether entanglement issues, rendering incomplete missions, as they cannot account for tether constraints. In simulation, REACT shows effective tether management while maintaining full coverage, achieving a shorter total mission time ($680$s vs. $855$s) despite longer inspection time due to its proactive entanglement avoidance that eliminates the need for extensive post-mission disentanglement. The real-world experiments confirm these findings, where the baseline CPP planner achieves only $80.9\%$ coverage before becoming physically stuck due to tether entanglement, while REACT completes the full mission with $95.6\%$ coverage through adaptive replanning. Additionally, REACT mitigates potential entanglement through reactive replanning, offering a more robust and reliable solution. REACT improves mission safety and integrity, making it ideal for tethered vehicle inspection scenarios where entanglement-awareness is critical.
@@ Acknowledgment
This research was supported by Innovation Fund Denmark (Grant No. 2040-00032B) and EIVA a/s. The authors thank Nicolas Dahn, Dr. Leif Christensen, and Tom Slawik from DFKI for their assistance with the experiments, and Adis Hodzic from EIVA for his technical support with the ROV control software. We also thank Dr. Louis Petit for his insights on the ropeRRT planning method and for open-sourcing his OMPL-based implementation and Dr. Olaya Álvarez-Tuñón for her valuable technical discussions.

### v2

@@ V-B Recovery path refinement
Once the recovery path $\mathbf{P}_{\text{recovery}}$ is generated through the disentanglement search, it undergoes a refinement process to generate a safe path $\mathbf{P}_{\text{safe}}$ with obstacle clearance $\delta_{obst}$ and improved smoothness prior to execution. This refinement is performed iteratively and consists of three consecutive operations: centroid offset, stochastic local perturbation, and polynomial smoothing. The resulting safe trajectory $\mathbf{P}_{\text{safe}}$ is validated to ensure collision-free execution.
The centroid offsetting step is designed to increase clearance from obstacles that may lie near the geometric center of the path. The geometric centroid of the path is calculated as the mean of all points, $\mathbf{c}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{p}_{i}$, where $N$ denotes the number of points in $\mathbf{P}_{\text{recovery}}$ and $\mathbf{p}_{i}$ represents the $i$-th point. Each point is then displaced along the normalized vector that points from the centroid to the point, $\mathbf{d}_{i}=(\mathbf{p}_{i}-\mathbf{c})/\|\mathbf{p}_{i}-\mathbf{c}\|$, producing the updated position $\mathbf{p}_{i}^{\prime}=\mathbf{p}_{i}+\delta_{obst}\mathbf{d}_{i}$. This operation effectively increases the clearance from the obstacles.
Following centroid offsetting, the path is further refined using stochastic local perturbations. Each intermediate point along the path (excluding the endpoints) is locally adjusted by sampling random directions uniformly on the surface of a sphere and displacing the point along the sampled direction by a small distance. The resulting candidate points are evaluated for collisions and only those that are collision-free are accepted. Multiple trials are allowed for each point to maximize the likelihood of finding a feasible local adjustment. This perturbation mechanism adjusts the path points to locally increase clearance from nearby obstacles, ensuring the trajectory remains safely away from potential collisions while preserving its overall structure. The final stage of refinement involves smoothing the path using a third-order polynomial fitting. The path is divided into overlapping segments and for each segment separate cubic polynomials are fitted to the x, y, and z coordinates of the points using a least squares approximation. Each point in the segment is then replaced by the corresponding evaluation of the polynomial functions, producing a smoother trajectory. The overlapped segments ensure continuity between the polynomial pieces and prevent abrupt changes in curvature, resulting in a dynamically feasible path suitable for the ROV. After completion of these operations, the refined path is subjected to collision verification. If any collisions are detected, the refinement process is repeated iteratively until a collision-free trajectory is obtained.
@@ VI Simulations Experiments
This section presents the simulation results for the proposed REACT method. The entire framework is implemented in C++ to ensure computational efficiency and is integrated with ROS to facilitate modularity and deployment on real-world robots. Furthermore, the open motion planning (OMPL) library [25] is utilized to compute the shortest paths.
@@ VI-A Simulation experimental setup
The path planner is implemented for a BlueROV2 underwater robot. An MPC approach accounts for model constraints and provides the optimal control input $\mathbf{u}^{ref}\in\mathbb{R}^{4}$, where $\mathbf{u}^{ref}=[F_{x},F_{y},F_{z},M_{z}]^{T}$ represents the forces in the $X$, $Y$, and $Z$ directions, and the moment about the $Z$-axis. The goal is to follow the desired reference trajectory $\mathbf{x}^{\text{ref}}\in\mathbb{R}^{4}$, where $\mathbf{x}^{\text{ref}}=[x^{\text{ref}},y^{\text{ref}},z^{\text{ref}},\psi^{\text{ref}}]^{T}$ contains the reference position $(x^{\text{ref}},y^{\text{ref}},z^{\text{ref}})$ and the reference yaw angle $\psi^{\text{ref}}$. For further details about the controller and model, the reader is referred to [4]. The entanglement-aware path planner provides $\mathbf{x}^{ref}$ in real-time for the MPC.
We perform a comparative analysis of the proposed REACT method and a baseline conventional CPP (FC-Planner [13]), which does not explicitly handle entanglement. The simulation setup uses a 1/10-scale pipe structure based on the model described in [13]. The simulated onboard camera has a $70$-degree field of view. The tether constraint is implemented by specifying a maximum allowable tether length of $L_{\text{max}}=10$m. Note that the resulting planned trajectories are geometrically identical to those in the full-scale environment, simply scaled down proportionally. The primary motivation for using the scaled model is to accelerate simulation runs while preserving the validity of the results. Coverage is computed in a manner similar to [5]. At each time step, the position and orientation of the camera are used to determine which triangles in the environment mesh are visible. A triangle is marked as visible if its centroid is within the inspection range, its surface normal faces the camera, and its projection lies within the camera’s field of view. Over time, the set of all uniquely observed triangles accumulates. The coverage at time $t$ is defined as the ratio of the number of unique visible triangles up to time $t$ to the total number of triangles in the environment.
@@ VI-B Simulations results
Table I: Coverage performance comparison between REACT and CPP baseline showing inspection time, recovery time, total mission duration, and final coverage.
@@
Table II: Comparison of tether constraint compliance.
@@
Figure 4: Comparison of coverage and tether length.
@@
The performance metrics for both planners are summarized in Table I and Table II, while Figs. 4 and 5 visualize the performance of the planners in terms of environmental coverage, tether length behavior, and final trajectory configurations. The evaluation focuses on comparing mission efficiency, constraint compliance, and safety aspects between the entanglement-aware REACT method and the conventional CPP..
@@
REACT demonstrates multiple instances of entanglement detection and resolution, as evidenced by the peaks in the tether length curve in Fig. 4(b) and the corresponding flat regions in the coverage curve in Fig. 4(a), where the ROV inspection progress stops to reroute and find paths without entanglement. This reactive replanning behavior directly results in a longer inspection time, as the system prioritizes tether safety over raw speed. In particular, the limit on the length of the tether is rarely exceeded by REACT, as shown in Table II and Fig. 4(b). Conversely, the CPP method severely violates this constraint, reaching 31.1631.16m for extended periods ($327.37$s), risking unrecoverable tether entanglement. The 3D trajectory visualizations in Fig. 5 further illustrate this difference, showing the entangled tether geometry resulting from CPP versus the non-entangled configuration achieved by REACT. Therefore, REACT demonstrates superior performance for safe tethered ROV inspection.
@@ VII Real-world Experiments
[…]
Figure 6: During the experiment, the ROV became entangled with the pipe, illustrating a typical CPP entanglement.
VII-A Real-world experimental results
@@
Real-world tests illustrate the need of REACT compared to classic CPP approach. Figure 7 shows the mission trajectories for both methods. As such, the baseline CPP planner failed to complete the inspection task. At $t=124$ s, the tether entangles the pipe structure, stopping the ROV from further progress. Hence, the overall coverage achieved is only $80.9\%$ (Table III).
@@
The proposed REACT planner completed the full mission, achieving $95.6\%$ coverage. The planner adapted to tether constraints throughout the mission. Unlike the simulation results, the soft tether length limit of the 88m was exceeded in real-world tests. As such, when no entanglement-free paths to the goal are available, REACT treats the length constraints as a soft limit rather than a hard limitation. However, when entanglement is detected and there is an entanglement-free path around the pipe, REACT finds a path that detangles the ROV from the structure. This capability is demonstrated at $t=96$s when REACT detected an imminent tether entanglement. The planner computed a new path that safely untangled the ROV from the pipe structure, allowing inspection to continue. The system only exceeded the soft tether limit when necessary for mission completion while prioritizing entanglement avoidance whenever possible.
@@
The baseline CPP planner, lacking entanglement-awareness, became permanently stuck (see Fig. 6). This demonstrates a key advantage of REACT that was not visible in the simulation study, where tether forces were not modeled. Although the REACT adds computation compared to the baseline, the maximum replanning time of 0.4822 s remained well within the onboard platform’s capabilities, allowing real-time operation. This overhead, while higher than the conventional CPP approach, was low enough to enable timely adaptation to tether entanglement events. These results demonstrate that standard planners without tether constraints fail even during relatively simple underwater 3D inspection tasks. While the baseline initially followed the planned path, it became entangled and subsequently immobilized midway through the mission. On the other hand, REACT’s ability to detect and avoid entanglement through replanning enabled a successful mission completion. These findings highlight the need for entanglement-aware planning in real underwater operations.
Table III: Coverage Performance in Real-world Trials. REACT takes a maximum of 0.4822 s (per re-planning step) during the inspection to compute the fully replanned path.
Planner Inspection Coverage (%) Computation overhead (s)
REACT Full 95.6 0.4822
CPP Failed (Entangled) 80.9 N/A
@@
Figure 7: Real-world ROV trajectories. Without REACT, the mission fails due to tether entanglement. With REACT, the ROV completes the mission successfully. At $t=96$s, the replanner is activated, allowing REACT to disentangle the ROV and continue the inspection. In contrast, at $t=124$s, the ROV using the conventional CPP becomes stuck.
@@ VIII Conclusion
We present REACT, a real-time, entanglement-aware path planning framework for tethered underwater vehicles that supports underwater asset inspection tasks. Through comparative simulation and real-world experiments, we demonstrate that conventional planners face tether entanglement issues, rendering incomplete missions, as they cannot account for tether constraints. In simulation, REACT shows effective tether management while maintaining full coverage, achieving a shorter total mission time ($680$s vs. $855$s) despite longer inspection time due to its proactive entanglement avoidance that eliminates the need for extensive post-mission disentanglement. The real-world experiments confirm these findings, where the baseline CPP planner achieves only $80.9\%$ coverage before becoming physically stuck due to tether entanglement, while REACT completes the full mission with $95.6\%$ coverage through adaptive replanning.
@@ Acknowledgment
This work was supported by Innovation Fund Denmark (Grant No. 2040-00032B) and EIVA a/s. The authors thank Nikolas, Tom, and Leif from DFKI for experimental support.